OpenStudy (anonymous):

art 1: Is the following equation true? (2 points) Part 2: Use complete sentences to explain the properties used in making your decision. (6 points) 3 log base 2 x + 1/2 log base 2 y – 3 log base 2 z = log base 2 x^3 sqrt y/z^3

OpenStudy (anonymous):

3 log base 2 (x) +1/2 log base 2 (y) -3 log base 2 (z)=log base 2 (x^3) + log base 2 (y^(1/2)) -log base 2 (z^3) remember, addition/subtraction of logs of same base and multiplication factors is done by: log(a)+log(b)=log(a*b), log(a)-log(b)=log(a/b) can you finish simplifying with that info?

OpenStudy (anonymous):

wow, yes thank you !!!!!!!!!'

OpenStudy (anonymous):

You know the rule that an exponent inside a log can be moved outside as a simple multiplier or the other way around?